[figure: 9.7 cm]

COMPLEX MAGNETIC STRUCTURE OF STRONGLY COUPLED Fe/Si MULTILAYERS

J. Dubowik, F. Stobiecki, B. Szymański^{a} Yu.V. Kudryavtsev^{b} A. Grabias and M. Kopcewicz^{c} aInstitute of Molecular Physics, PASci, Smoluchowskiego 17, 60–179 Poznań bInstitute of Metal Physics, NASci of Ukraine, Kiev-142, Ukraine cInstitute of Electronic Materials Technology, Wólczyńska 133, 01–919 Warsaw

Fe/Si multilayers exhibiting a strong antiferromagnetic (AF) coupling for a nominal Si spacer thickness 10 < t_{Si} < 17 Å have been investigated with various magnetic techniques. The in-plane magnetization measurements revealed maximal coupling strength for t_{Si}=11–13 Å with a saturation field H_{sat} ≈ 10–12 kOe. A biquadratic contribution of the same order as the bilinear term in exchange coupling was observed. Conversionelectron Mössbauer spectroscopy measurements suggest a complex nature of interface magnetic structure due to diffusion of Si into Fe. As a result, only 20–30 % of Fe reveals α–Fe structure, about 60 % constitutes Fe rich Fe-Si alloy and about 10 % of Fe is nonmagnetic. Ferromagnetic resonance spectra taken at 9.1 and 33.6 GHz consist of the fundamental mode and higher order modes which lie both on the low and the high magnetic field side of the fundamental mode for t_{Si} < 15 Å. This is attributed to a mixed F/AF coupling resulting in the apparent strong biquadratic term. Dispersion relations for AF coupled bilayer structures, frequently used for analysing FMR in multilayers, are shown to be not applicable to Fe/Si multilayers.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : J. Dubowik

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences Smoluchowskiego 17 60-179 Poznan

Email address : dubowik@ifmpan.poznan.pl